OpenStudy (anonymous):

Nisha can solve a set of 5 math problems in 12 minutes. At this rate, how long will it take her to solve 20 sets of 7 math problems? F 58 minutes G 5 hours 16 minutes H 48 minutes J 5 hours 36 minutes

OpenStudy (anonymous):

it is letter J.

OpenStudy (anonymous):

just follow these steps. 5/12=7/x x there is the time for a set do the cross multiplication. 5x=84 x=16.8, but there are 20 sets so, all we have to do is multiply this by 20. x=16.8(20)= 336 minutes 336 minutes is equal to 5 hours and 36 minutes.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you are welcome. can you become my fan?